### v1

@@ 0.C.2 A Better Lower Bound for tbn(G)
Before introducing the next intermediate result, which is Lemma 10, we need to define more concepts.
Let $C$ be a demanding cycle and let $P$ be a contour path of $C$. Path $P$ is interesting if it is a sub-path of some contour paths in all non-demanding cycles of $G$ containing $C$ and sharing some edges with $C$. See Fig. 14(a) and consider $C$, whose legs are $\{e_{1},e_{2},e_{3}\}$; $C^{\prime}$, whose legs are $\{e_{1},e_{4},e_{5}\}$, and $C^{\prime\prime}$, whose legs are $\{e_{6},e_{7},e_{8}\}$. Cycle $C^{\prime\prime}$ contains $C^{\prime}$, and $C^{\prime}$ contains $C$. Cycle $C^{\prime}$ is demanding and its contour path $P$ is interesting, as it is part of both $C^{\prime}$ and $C^{\prime\prime}$. On the other hand, $\tilde{P}$ is not interesting, as it is part of $C^{\prime}$ but not of $C^{\prime\prime}$.
Lemma 9
@@
A demanding cycle $C\in D(G)$ is expensive if and only if the following two properties hold (refer to Fig. 14(a) for an example).
@@
See for example Fig. 13(a), where $C_{5}$ and $C_{9}$ are expensive. We have that $C_{4}$ and $C_{2}$ are the the non-demanding cycle sharing exactly one edge with $C_{5}$ and $C_{9}$ and not sharing edges with other demanding cycles.
Let $D_{\exp}(G)\subseteq D(G)$ be the set of the expensive cycles of $G$. The following lemma refines the lower bound defined in Corollary 1.
Lemma 10
@@
Proof
Let $\mathcal{C}$ be an unbent collection with $b=\operatorname{tbn}(G)$ dummies. We show that $\mathcal{C}$ has at least $2|D(G)|+|D_{\exp}(G)|+8-\min\{8,2|D_{\operatorname{ext}}(G)|\}$ dummies. If $|D_{\exp}(G)|=0$ the lemma follows by Corollary 1. Suppose $|D_{\exp}(G)|>0$. Consider an expensive demanding cycle $C$. Let $e$ be the edge forming the only interesting contour path of $C$. Let $C^{\prime}$ be the non-demanding cycle containing $C$, sharing edges with the interesting contour path of $C$ (i.e., $e$ is along the boundary of $C$ and $C^{\prime}$), and not containing other non-demanding cycles with this property (see Property 1). Cycle $C^{\prime}$ does not share the dummy with $C$ in at least one of the representations, otherwise $e$ (which is the only edge they share) would not have a representation where it is not subdivided. It follows that, in such representation, in order to make both $C$ and $C^{\prime}$ good cycles, which is necessary by Lemma 4, we need two dummies, one for $C$ and one for $C^{\prime}$. In all the other representations, we need one dummy, that is shared by $C$ and $C^{\prime}$ and placed along $e$. Notice that expensive cycles are contained in $D(G)$ and consequently they are vertex disjoint. Hence, the above consideration hold for every expensive cycle independently. It follows that, if $|\mathcal{C}|=2$ (since $G$ is triconnected, this is the minimum possible cardinality of $\mathcal{C}$), $\mathcal{C}$ has at least $|D_{\exp}(G)|$ dummies in addition to the ones of Corollary 1. Hence, the lemma follows.
We now refine the lower bound of Lemma 10, introducing Lemma 5. Before that we show the following lemma, that states a relationship between the contour paths of the cycles in $D(G)$ and the external face $f_{ext}$. In particular, a contour path $P$, that is a contour path of a cycle $C\in D(G)$ and sharing edges with $f_{\operatorname{ext}}$ ($f_{\operatorname{ext}}$ is a leg face of $C$), is always a sub-path of a contour path of a non-demanding cycle $C^{\prime}$ containing $C$ (if $C^{\prime}$ exists). Consequently, $P$ is always interesting. The same holds with respect to faces incident to $f_{\operatorname{ext}}$, when $f_{\operatorname{ext}}$ is not a leg face of $C$. More precisely, we have the following.
Lemma 11 ([5])
@@
A cycle $C\in D_{\operatorname{ext}}(G)$ is short if and only if it has (i) exactly one edge $e$ incident to $f_{\operatorname{ext}}$ and (ii) two or more interesting contour paths. Observe that $e$ is a contour path (in particular, a contour path consisting of exactly one edge) of $C$ and that it is interesting by Lemma 11. We denote by $D_{\operatorname{short}}(G)$ the set of the short demanding cycles. Observe that $D_{\operatorname{short}}(G)\cap D_{\exp}(G)=\emptyset$, as (ii) contradicts 1. For an example, see Fig. 14(b), where there are two 3-legged cycles $C$ (with legs $\{e_{1},e_{2},e_{3}\}$) and $C^{\prime}$ (with legs $\{e_{1},e_{4},e_{5}\}$). Note that $C$ is demanding and that the contour paths $P$ and $\tilde{P}$ of $C$ are interesting, as they are part of $C^{\prime}$. Also, $C$ is short, as it has two interesting contour paths ($P$ and $\tilde{P}$) and since $P$ is composed of one edge and incident to $f_{\operatorname{ext}}$.
@@
Thanks to the concept of short demanding cycle we can refine the lower bound of Lemma 10 with the following lemma. Also, the following lower bound is tight, as we are going to prove in the next section.
See 5
Proof
@@
Let $\mathcal{C}$ be an unbent collection with $b=\operatorname{tbn}(G)$ dummies. If $2|D_{\operatorname{ext}}(G)|-|D_{\operatorname{short}}(G)|\geq 8$ or if $|D_{\operatorname{short}}(G)|=\emptyset$, then the bounds in Lemma 10 and Lemma 5 coincide. Hence, suppose $2|D_{\operatorname{ext}}(G)|-|D_{\operatorname{short}}(G)|\leq 7$ and $|D_{\operatorname{short}}(G)|\geq 1$. We have to show that in this case there are $|D_{\operatorname{short}}(G)|$ additive dummies in $\mathcal{C}$ with respect to the ones of Lemma 10. Consider $C\in D_{\operatorname{short}}(G)$ and let $e$ be the edge of $C$ incident to $f_{\operatorname{ext}}$, forming a contour path of $C$. Recall that, by Lemma 11, such contour path is interesting. Hence, placing a dummy on $e$ is optimal, as in this case such dummy place the role of: A dummy for $C$ and all the non-demanding sharing an edge with $C$; one of the 4 dummies in the external face. On the other hand, in our setting it is not possible that there is a dummy on $e$ in all the representations of $\mathcal{C}$. Let $H\in\mathcal{C}$ be the representation where there is no dummy in the edge $e$ of $C$ incident to $f_{\operatorname{ext}}$. In such representation, we have to place a dummy in an edge of $C$ belonging to an interesting contour path of $C$ (which exists by definition of short cycles) not incident to $f_{\operatorname{ext}}$ and in another edge of the external face of $G$, instead of placing just on $e$. In all the others representation, just a dummy placed on $e$ suffices. Since $D_{\operatorname{ext}}(G)\leq 7$ and since $C\in D_{\operatorname{ext}}(G)$, the cost with respect to the optimum increases of at least 1 in $H$. This consideration holds for each element of $D_{\operatorname{short}}(G)$. If $|\mathcal{C}|=2$ (which is the minimum we can have) $\mathcal{C}$ has at least $|D_{\operatorname{short}}(G)|$ dummies in addition to the ones of Lemma 10 and, in particular, the lemma follows.
@@
Figure 15: An unbent collection of the graph $G$ depicted in Fig. 13(a) with dummies (blue circles) whose number matches the lower bound of Lemma 5.
@@
In Section 0.C.3 we are going to show that the lower bound of Lemma 5 is tight. Fig. 15 shows an unbent collection of $G$ depicted in Fig. 13(a), where dummies are represented as disks. As observed before, for such $G$ we have $D(G)=\{C_{3},C_{5},C_{8}.C_{9}\}$ and $D_{\operatorname{ext}}(G)=\{C_{3},C_{8}.C_{9}\}$. We also have $D_{\exp}(G)=\{C_{5},C_{9}\}$ and $D_{\operatorname{short}}(G)=\{C_{8}\}$. We have that such collection has overall $\operatorname{tbn}(G)=2|D(G)|+|D_{\exp}(G)|+8-\min\{8,2|D_{\operatorname{ext}}(G)|-|D_{\operatorname{short}}(G)|\}=8+2+8-5=13$ dummies.
0.C.3 Computing a Collection with tbn(G) Bends
@@
In this section we show that the lower bound expressed in Lemma 5 is tight and we prove Theorem 4.1. A 3-legged cycle is maximal if it is not contained in another 3-legged cycle. We use the following lemma to prove Lemma 13, which is a key ingredient for Theorem 4.1.
Lemma 12 ([5, 13])
@@
Proof
By Lemma 12, the set $D(G)$ can be computed in $O(n)$ time and also their interesting contour paths. Since they are not intersecting, they are independent. Hence, we can traverse all their edges. Whenever we find a contour path that is composed by an edge, we verify if it is an interesting contour path and if it is the only interesting contour path of the cycle. In this case, the cycle is expensive by definition and we add it to $D_{\exp}(G)$, otherwise it is not and we do nothing. Concerning $D_{\operatorname{short}}(G)$, we visit each cycle of $D(G)$ and we count the interesting contour paths of each one of them. For the ones with 2 or more interesting contour paths, we test if there exists one incident to $f_{\operatorname{ext}}$ and if such contour path is composed by one edge. If it exists, we add the cycle to $D_{\operatorname{short}}(G)$.
We now turn to Lemma 17, which is another key result to prove Theorem 4.1. Before that, we focus on the concept of a genealogical tree $T(C)$ for a 3-legged cycle $C$. In particular, we enrich $T(C)$ with more information, as follows.
Lemma 14 ([13])
@@
We say that a non-demanding cycle $C^{\prime}$ is critical if, for every demanding cycle $C^{\prime\prime}$ it contains, $C^{\prime}$ shares exactly one edge with $C^{\prime\prime}$ and such edge corresponds to an interesting contour path of $C^{\prime\prime}$. Cycles $C_{5}$, $C_{6}$, and $C_{9}$ in Fig. 16(a) are critical.
@@
See for example Fig. 16. Fig. 16(a) depicts $G_{C}$ and Fig. 16(b) represents $T^{+}(G)$. Observe that 1 holds, as $\alpha(C)\not=\emptyset$ for any 3-legged cycle; also, we have that $\alpha(C_{5})=C_{4}$ and $\beta(C_{5})=\emptyset$ and $C_{4}$ is an expensive demanding cycle. Observe also that 1 holds: For example, we have that $\alpha(C_{9})=C_{2}$, $\beta(C_{6})=C_{1}$ and, consequently, $\beta(C_{9})=C_{1}$.
Lemma 15
@@
Proof
We attach to the three degree-2 vertices in the external cycle of $G_{C}$ a 3-cycle so that $G_{C}$ is now cubic. Then, we compute $T(C)$ by Lemma 14, and $D(G_{C})$ and the relative interesting contour paths by Lemma 12. We then remove the external 3-cycle.
@@
By Lemma 14 two cycles are never intersecting in $T(C)$. Hence, every edge of $G_{C}$ is visited $O(1)$ times thanks to the pointers that we add to the legs of the non-demanding children of every currently visited cycle. It follows that the enriched genealogical tree $T^{+}(C)$ can be computed in $O(n_{C})$ time.
Lemma 16
@@
Proof
See Fig. 14(a) for an example of the statement of the lemma, where $P$ is an interesting path, $e_{2}$ is leg of the demanding cycle $C$, and $e_{2}$ is also and edge of $C^{\prime}$ and of $C^{\prime\prime}$. Let $l_{1}$ and $l_{2}$ be the two legs of $C$ incident to vertices of $P$. Since $C^{\prime}$ and $C^{\prime\prime}$ are non-intersecting, it is not possible that we both have that: $l_{1}$ is an edge of $C^{\prime}$ and not an edge of $C^{\prime\prime}$; and that $l_{2}$ is an edge of $C^{\prime\prime}$ and not an edge of $C^{\prime}$.
@@
Proof
We compute the enriched genealogical tree $T^{+}(C)$ in $O(n_{C})$ time by Lemma 15. We also compute $D(G_{C})$, $D_{\exp}(G_{C})$, and their interesting contour paths by means of Lemma 12 and Lemma 13. We consider two copies of $G_{C}$, namely $G^{1}_{C}$ and as $G^{2}_{C}$ and place dummies along their edges (i.e., we compute $\mathcal{C}$ such that $|\mathcal{C}|=2$). The proof consists of two parts: In Part 1 we describe an $O(n_{C})$-time procedure that subdivides edges of $G^{1}_{C}$ and $G^{2}_{C}$ with $2|D(G_{C})|+|D_{\exp}(G_{C})|$ dummies so that neither $G^{1}_{C}$ nor $G^{2}_{C}$ has a bad $3$-legged cycle. In this part we prove both (i) and (ii). In Part 2 we prove that at the end of this subdivision procedure neither $G^{1}_{C}$ nor $G^{2}_{C}$ contains a bad cycle (observe that in this case a bad cycle could be the external cycle or a 2-legged cycle generated by the placement of the dummy vertices, see for example $C_{1}$ and $C_{2}$ in Fig. 12). Finally, we discuss the computational time.
Part 1. We first focus on $(i)$. The strategy is counting the number of dummies that we place in both $G^{1}_{C}$ and $G^{2}_{C}$ and that exceed the number $2|D(G_{C})|$ (for example, if analyzing a cycle in $D(G_{C})$ we place 3 dummies overall in the collection, we are exceeding 1). Observe first that $G_{C}$ has no intersecting cycles by Lemma 14. Hence, we can focus on pairwise non-intersecting cycles. Consider each demanding cycle in $T^{+}(C)$ with at least two edges belonging to an interesting contour path. We consider two edges of it following this rule:
@@
When we consider a critical cycle $C^{\prime}$, we denote by $e_{x}$ the edge forming the interesting contour path of $\alpha(C^{\prime})$ contained in $C^{\prime}$. We define $e_{y}$ similarly, if $\beta(C^{\prime})\not=\emptyset$. During the procedure, we implicitly assume that if we subdivide $e_{x}$ in $G^{a}_{C}$ ($a\in\{1,2\}$), we subdivide another edge of a contour path of $\alpha(C^{\prime})$ that is not interesting in $G^{b}_{C}$ ($b\in\{1,2\}\setminus a$), so that after this placement $\alpha(C^{\prime})$ is not a bad cycle in both $G^{1}_{C}$ and $G^{2}_{C}$ (in some case we explicitly say how to perform also this subdivision, if this is helpful for the description of the algorithm). We do that also for $\beta(C^{\prime})$ when it is not equal to $\emptyset$. During the procedure, we maintain three invariants:
@@
Consider the case $i=1$. If $\beta(C^{\prime})\not=\emptyset$, we subdivide $e_{x}$ in $G^{1}_{C}$ and $e_{y}$ in $G^{2}_{C}$ and we set $count_{1}=count_{0}=0$. Hence, 1 holds. Otherwise, we subdivide $e_{x}$ in $G^{1}_{C}$, while for $G^{2}_{C}$ we consider the leg $l$ of $\alpha(C^{\prime})$ whose existence is stated in Lemma 16 and any edge $e_{\tilde{x}}$ of $\alpha(C^{\prime})$ different from $e_{x}$. We subdivide both of them and, by Lemma 16, none of the non-demanding cycles sharing $e_{x}$ with $\alpha(C^{\prime})$ are bad in both $G^{1}_{C}$ and $G^{2}_{C}$. By the same lemma invariant 3 holds. In this case $count_{1}=count_{0}+1=1$, as we placed two dummies instead of one considering $\alpha(C^{\prime})$ in $G^{2}_{C}$ (one dummy in $e_{\tilde{x}}$ and one in $l$). Observe that in this case 2 trivially holds as this is the first step.
@@
Suppose by contradiction that $e_{x}$ and $e_{y}$ are both subdivided in one between $G^{1}_{C}$ and $G^{2}_{C}$. By 1, we have that we did not subdivide them in the same graph in step $j\in[1,j-1]$. Hence, they were subdivided in two different steps, while analyzing two different cycles $C_{p}$ and $C_{p^{\prime}}$. As they share at least one edge and by Lemma 14, either $C_{p}$ contains $C_{p^{\prime}}$ or vice versa. Suppose $C_{p^{\prime}}$ contains $C_{p}$. By 2, if $C_{p^{\prime}}$ was associated to $e_{x}$ or $e_{y}$, the same holds for $C_{p}$. By 2, if one between $e_{x}$ and $e_{y}$ was previously subdivided in $G^{1}_{C}$ by $C_{p^{\prime}}$, $C_{p}$ did not subdivide the other one in $G^{1}_{C}$. A contradiction. Hence, $e_{x}$ and $e_{y}$ are not both subdivided in one between $G^{1}_{C}$ and $G^{2}_{C}$.

### v2

@@ C.2 A Better Lower Bound for tbn(G)
Before introducing the next intermediate result, which is Lemma C.6, we need to define more concepts.
Let $C$ be a demanding cycle and let $P$ be a contour path of $C$. Path $P$ is interesting if it is a sub-path of some contour paths in all non-demanding cycles of $G$ containing $C$ and sharing some edges with $C$. See Figure 14(a) and consider $C$, whose legs are $\{e_{1},e_{2},e_{3}\}$; $C^{\prime}$, whose legs are $\{e_{1},e_{4},e_{5}\}$, and $C^{\prime\prime}$, whose legs are $\{e_{6},e_{7},e_{8}\}$. Cycle $C^{\prime\prime}$ contains $C^{\prime}$, and $C^{\prime}$ contains $C$. Cycle $C^{\prime}$ is demanding and its contour path $P$ is interesting, as it is part of both $C^{\prime}$ and $C^{\prime\prime}$. On the other hand, $\tilde{P}$ is not interesting, as it is part of $C^{\prime}$ but not of $C^{\prime\prime}$.
Lemma C.5.
@@
A demanding cycle $C\in D(G)$ is expensive if and only if the following two properties hold (refer to Figure 14(a) for an example).
@@
See for example Figure 13(a), where $C_{5}$ and $C_{9}$ are expensive. We have that $C_{4}$ and $C_{2}$ are the the non-demanding cycle sharing exactly one edge with $C_{5}$ and $C_{9}$ and not sharing edges with other demanding cycles.
Let $D_{\exp}(G)\subseteq D(G)$ be the set of the expensive cycles of $G$. The following lemma refines the lower bound defined in Corollary C.4.
Lemma C.6.
@@
Proof C.7.
Let $\mathcal{C}$ be an unbent collection with $b=\operatorname{tbn}(G)$ dummies. We show that $\mathcal{C}$ has at least $2|D(G)|+|D_{\exp}(G)|+8-\min\{8,2|D_{\operatorname{ext}}(G)|\}$ dummies. If $|D_{\exp}(G)|=0$ the lemma follows by Corollary C.4. Suppose $|D_{\exp}(G)|>0$. Consider an expensive demanding cycle $C$. Let $e$ be the edge forming the only interesting contour path of $C$. Let $C^{\prime}$ be the non-demanding cycle containing $C$, sharing edges with the interesting contour path of $C$ (i.e., $e$ is along the boundary of $C$ and $C^{\prime}$), and not containing other non-demanding cycles with this property (see Property 1). Cycle $C^{\prime}$ does not share the dummy with $C$ in at least one of the representations, otherwise $e$ (which is the only edge they share) would not have a representation where it is not subdivided. It follows that, in such representation, in order to make both $C$ and $C^{\prime}$ good cycles, which is necessary by Lemma 4.1, we need two dummies, one for $C$ and one for $C^{\prime}$. In all the other representations, we need one dummy, that is shared by $C$ and $C^{\prime}$ and placed along $e$. Notice that expensive cycles are contained in $D(G)$ and consequently they are vertex disjoint. Hence, the above consideration hold for every expensive cycle independently. It follows that, if $|\mathcal{C}|=2$ (since $G$ is triconnected, this is the minimum possible cardinality of $\mathcal{C}$), $\mathcal{C}$ has at least $|D_{\exp}(G)|$ dummies in addition to the ones of Corollary C.4. Hence, the lemma follows.
We now refine the lower bound of Lemma C.6, introducing Lemma 4.2. Before that we show the following lemma, that states a relationship between the contour paths of the cycles in $D(G)$ and the external face $f_{ext}$. In particular, a contour path $P$, that is a contour path of a cycle $C\in D(G)$ and sharing edges with $f_{\operatorname{ext}}$ ($f_{\operatorname{ext}}$ is a leg face of $C$), is always a sub-path of a contour path of a non-demanding cycle $C^{\prime}$ containing $C$ (if $C^{\prime}$ exists). Consequently, $P$ is always interesting. The same holds with respect to faces incident to $f_{\operatorname{ext}}$, when $f_{\operatorname{ext}}$ is not a leg face of $C$. More precisely, we have the following.
Lemma C.8 ([6]).
@@
A cycle $C\in D_{\operatorname{ext}}(G)$ is short if and only if it has (i) exactly one edge $e$ incident to $f_{\operatorname{ext}}$ and (ii) two or more interesting contour paths. Observe that $e$ is a contour path (in particular, a contour path consisting of exactly one edge) of $C$ and that it is interesting by Lemma C.8. Let $D_{\operatorname{short}}(G)$ be the set of the short demanding cycles. Observe that $D_{\operatorname{short}}(G)\cap D_{\exp}(G)=\emptyset$, as (ii) contradicts 1. For an example, see Figure 14(b), where there are two 3-legged cycles $C$ (with legs $\{e_{1},e_{2},e_{3}\}$) and $C^{\prime}$ (with legs $\{e_{1},e_{4},e_{5}\}$). Note that $C$ is demanding and that the contour paths $P$ and $\tilde{P}$ of $C$ are interesting, as they are part of $C^{\prime}$. Also, $C$ is short, as it has two interesting contour paths ($P$ and $\tilde{P}$) and since $P$ is composed of one edge and incident to $f_{\operatorname{ext}}$.
@@
Thanks to the concept of short demanding cycle we can refine the lower bound of Lemma C.6 with the following lemma. Also, the following lower bound is tight, as we are going to prove in the next section.
See 4.2
Proof C.9.
@@
Let $\mathcal{C}$ be an unbent collection with $b=\operatorname{tbn}(G)$ dummies. If $2|D_{\operatorname{ext}}(G)|-|D_{\operatorname{short}}(G)|\geq 8$ or if $|D_{\operatorname{short}}(G)|=\emptyset$, then the bounds in Lemma C.6 and Lemma 4.2 coincide. Hence, suppose $2|D_{\operatorname{ext}}(G)|-|D_{\operatorname{short}}(G)|\leq 7$ and $|D_{\operatorname{short}}(G)|\geq 1$. We have to show that in this case there are $|D_{\operatorname{short}}(G)|$ additive dummies in $\mathcal{C}$ with respect to the ones of Lemma C.6. Consider $C\in D_{\operatorname{short}}(G)$ and let $e$ be the edge of $C$ incident to $f_{\operatorname{ext}}$, forming a contour path of $C$. Recall that, by Lemma C.8, such contour path is interesting. Hence, placing a dummy on $e$ is optimal, as in this case such dummy place the role of: A dummy for $C$ and all the non-demanding sharing an edge with $C$; one of the 4 dummies in the external face. On the other hand, in our setting it is not possible that there is a dummy on $e$ in all the representations of $\mathcal{C}$. Let $H\in\mathcal{C}$ be the representation where there is no dummy in the edge $e$ of $C$ incident to $f_{\operatorname{ext}}$. In such representation, we have to place a dummy in an edge of $C$ belonging to an interesting contour path of $C$ (which exists by definition of short cycles) not incident to $f_{\operatorname{ext}}$ and in another edge of the external face of $G$, instead of placing just on $e$. In all the others representation, just a dummy placed on $e$ suffices. Since $D_{\operatorname{ext}}(G)\leq 7$ and since $C\in D_{\operatorname{ext}}(G)$, the cost with respect to the optimum increases of at least 1 in $H$. This consideration holds for each element of $D_{\operatorname{short}}(G)$. If $|\mathcal{C}|=2$ (which is the minimum we can have) $\mathcal{C}$ has at least $|D_{\operatorname{short}}(G)|$ dummies in addition to the ones of Lemma C.6 and, in particular, the lemma follows.
@@
Figure 15: An unbent collection of the graph $G$ depicted in Figure 13(a) with dummies (blue circles) whose number matches the lower bound of Lemma 4.2.
@@
In Section C.3 we are going to show that the lower bound of Lemma 4.2 is tight. Figure 15 shows an unbent collection of $G$ depicted in Figure 13(a), where dummies are represented as disks. As observed before, for such $G$ we have $D(G)=\{C_{3},C_{5},C_{8}.C_{9}\}$ and $D_{\operatorname{ext}}(G)=\{C_{3},C_{8}.C_{9}\}$. We also have $D_{\exp}(G)=\{C_{5},C_{9}\}$ and $D_{\operatorname{short}}(G)=\{C_{8}\}$. We have that such collection has overall $\operatorname{tbn}(G)=2|D(G)|+|D_{\exp}(G)|+8-\min\{8,2|D_{\operatorname{ext}}(G)|-|D_{\operatorname{short}}(G)|\}=8+2+8-5=13$ dummies.
C.3 Computing a Collection with tbn(G) Bends
@@
In this section we show that the lower bound expressed in Lemma 4.2 is tight and we prove Theorem 4.3. A 3-legged cycle is maximal if it is not contained in another 3-legged cycle. We use the following lemma to prove Lemma C.11, which is a key ingredient for Theorem 4.3.
Lemma C.10 ([6, 13]).
@@
Proof C.12.
By Lemma C.10, the set $D(G)$ can be computed in $O(n)$ time and also their interesting contour paths. Since they are not intersecting, they are independent. Hence, we can traverse all their edges. Whenever we find a contour path that is composed of an edge, we verify if it is an interesting contour path and if it is the only interesting contour path of the cycle. In this case, the cycle is expensive by definition and we add it to $D_{\exp}(G)$, otherwise it is not and we do nothing. Concerning $D_{\operatorname{short}}(G)$, we visit each cycle of $D(G)$ and we count the interesting contour paths of each one of them. For the ones with 2 or more interesting contour paths, we test if there exists one incident to $f_{\operatorname{ext}}$ and if such contour path is composed of one edge. If it exists, we add the cycle to $D_{\operatorname{short}}(G)$.
We now turn to Lemma C.18, which is another key result to prove Theorem 4.3. Before that, we focus on the concept of a genealogical tree $T(C)$ for a 3-legged cycle $C$. In particular, we enrich $T(C)$ with more information, as follows.
Lemma C.13 ([13]).
@@
We say that a non-demanding cycle $C^{\prime}$ is critical if, for every demanding cycle $C^{\prime\prime}$ it contains, $C^{\prime}$ shares exactly one edge with $C^{\prime\prime}$ and such edge corresponds to an interesting contour path of $C^{\prime\prime}$. Cycles $C_{5}$, $C_{6}$, and $C_{9}$ in Figure 16(a) are critical.
@@
See for example Figure 16. Figure 16(a) depicts $G_{C}$ and Figure 16(b) represents $T^{+}(G)$. Observe that 1 holds, as $\alpha(C)\not=\emptyset$ for any 3-legged cycle; also, we have that $\alpha(C_{5})=C_{4}$ and $\beta(C_{5})=\emptyset$ and $C_{4}$ is an expensive demanding cycle. Observe also that 1 holds: For example, we have that $\alpha(C_{9})=C_{2}$, $\beta(C_{6})=C_{1}$ and, consequently, $\beta(C_{9})=C_{1}$.
Lemma C.14.
@@
Proof C.15.
We attach to the three degree-2 vertices in the external cycle of $G_{C}$ a 3-cycle so that $G_{C}$ is now cubic. Then, we compute $T(C)$ by Lemma C.13, and $D(G_{C})$ and the relative interesting contour paths by Lemma C.10. We then remove the external 3-cycle.
@@
By Lemma C.13 two cycles are never intersecting in $T(C)$. Hence, every edge of $G_{C}$ is visited $O(1)$ times thanks to the pointers that we add to the legs of the non-demanding children of every currently visited cycle. It follows that the enriched genealogical tree $T^{+}(C)$ can be computed in $O(n_{C})$ time.
Lemma C.16.
@@
Proof C.17.
See Figure 14(a) for an example of the statement of the lemma, where $P$ is an interesting path, $e_{2}$ is leg of the demanding cycle $C$, and $e_{2}$ is also and edge of $C^{\prime}$ and of $C^{\prime\prime}$. Let $l_{1}$ and $l_{2}$ be the two legs of $C$ incident to vertices of $P$. Since $C^{\prime}$ and $C^{\prime\prime}$ are non-intersecting, it is not possible that we both have that: $l_{1}$ is an edge of $C^{\prime}$ and not an edge of $C^{\prime\prime}$; and that $l_{2}$ is an edge of $C^{\prime\prime}$ and not an edge of $C^{\prime}$.
@@
Proof C.19.
We compute the enriched genealogical tree $T^{+}(C)$ in $O(n_{C})$ time by Lemma C.14. We also compute $D(G_{C})$, $D_{\exp}(G_{C})$, and their interesting contour paths by means of Lemma C.10 and Lemma C.11. We consider two copies of $G_{C}$, namely $G^{1}_{C}$ and as $G^{2}_{C}$ and place dummies along their edges (i.e., we compute $\mathcal{C}$ such that $|\mathcal{C}|=2$). The proof consists of two parts: In Part 1 we describe an $O(n_{C})$-time procedure that subdivides edges of $G^{1}_{C}$ and $G^{2}_{C}$ with $2|D(G_{C})|+|D_{\exp}(G_{C})|$ dummies so that neither $G^{1}_{C}$ nor $G^{2}_{C}$ has a bad $3$-legged cycle. In this part we prove both (i) and (ii). In Part 2 we prove that at the end of this subdivision procedure neither $G^{1}_{C}$ nor $G^{2}_{C}$ contains a bad cycle (observe that in this case a bad cycle could be the external cycle or a 2-legged cycle generated by the placement of the dummy vertices, see for example $C_{1}$ and $C_{2}$ in Figure 12). Finally, we discuss the computational time.
Part 1. We first focus on $(i)$. The strategy is counting the number of dummies that we place in both $G^{1}_{C}$ and $G^{2}_{C}$ and that exceed the number $2|D(G_{C})|$ (for example, if analyzing a cycle in $D(G_{C})$ we place 3 dummies overall in the collection, we are exceeding 1). Observe first that $G_{C}$ has no intersecting cycles by Lemma C.13. Hence, we can focus on pairwise non-intersecting cycles. Consider each demanding cycle in $T^{+}(C)$ with at least two edges belonging to an interesting contour path. We consider two edges of it following this rule:
@@
When we consider a critical cycle $C^{\prime}$, let $e_{x}$ be the edge forming the interesting contour path of $\alpha(C^{\prime})$ contained in $C^{\prime}$. If $\beta(C^{\prime})\not=\emptyset$, then we define $e_{y}$ similarly. During the procedure, we implicitly assume that if we subdivide $e_{x}$ in $G^{a}_{C}$ ($a\in\{1,2\}$), we subdivide another edge of a contour path of $\alpha(C^{\prime})$ that is not interesting in $G^{b}_{C}$ ($b\in\{1,2\}\setminus a$), so that after this placement $\alpha(C^{\prime})$ is not a bad cycle in both $G^{1}_{C}$ and $G^{2}_{C}$ (in some case we explicitly say how to perform also this subdivision, if this is helpful for the description of the algorithm). We do that also for $\beta(C^{\prime})$ when it is not equal to $\emptyset$. During the procedure, we maintain three invariants:
@@
Consider the case $i=1$. If $\beta(C^{\prime})\not=\emptyset$, we subdivide $e_{x}$ in $G^{1}_{C}$ and $e_{y}$ in $G^{2}_{C}$ and we set $count_{1}=count_{0}=0$. Hence, 1 holds. Otherwise, we subdivide $e_{x}$ in $G^{1}_{C}$, while for $G^{2}_{C}$ we consider the leg $l$ of $\alpha(C^{\prime})$ whose existence is stated in Lemma C.16 and any edge $e_{\tilde{x}}$ of $\alpha(C^{\prime})$ different from $e_{x}$. We subdivide both of them and, by Lemma C.16, none of the non-demanding cycles sharing $e_{x}$ with $\alpha(C^{\prime})$ are bad in both $G^{1}_{C}$ and $G^{2}_{C}$. By the same lemma invariant 3 holds. In this case $count_{1}=count_{0}+1=1$, as we placed two dummies instead of one considering $\alpha(C^{\prime})$ in $G^{2}_{C}$ (one dummy in $e_{\tilde{x}}$ and one in $l$). Observe that in this case 2 trivially holds as this is the first step.
@@
Suppose by contradiction that $e_{x}$ and $e_{y}$ are both subdivided in one between $G^{1}_{C}$ and $G^{2}_{C}$. By 1, we have that we did not subdivide them in the same graph in step $j\in[1,j-1]$. Hence, they were subdivided in two different steps, while analyzing two different cycles $C_{p}$ and $C_{p^{\prime}}$. As they share at least one edge and by Lemma C.13, either $C_{p}$ contains $C_{p^{\prime}}$ or vice versa. Suppose $C_{p^{\prime}}$ contains $C_{p}$. By 2, if $C_{p^{\prime}}$ was associated to $e_{x}$ or $e_{y}$, the same holds for $C_{p}$. By 2, if one between $e_{x}$ and $e_{y}$ was previously subdivided in $G^{1}_{C}$ by $C_{p^{\prime}}$, $C_{p}$ did not subdivide the other one in $G^{1}_{C}$. A contradiction. Hence, $e_{x}$ and $e_{y}$ are not both subdivided in one between $G^{1}_{C}$ and $G^{2}_{C}$.